The net amount of electricity produced in the U.S. (in billions of kilowatt hours) is approximated by function f(x)=68.6x+3009, where x=1 corresponds to 1996.
Using the model, find the year in which energy consumption firest exceeds 4300 billion kilowatt hours.

Answer by checkley77(12844)

f(x)=68.6x+3009
4,300=68.6x+3,009
68.6x=4,300-3,009
68.6x=1,291
x=1,251/68.6
x=18.8
1996+18.8=2014.8 is the year with 4,300 billion KWH.